Aphrodite Redux : Edna Pontellier’s Dilemma

in The Awakening by Kate Chopin

BY

Anne FOATA

It must be a shared experience in our response to a work of art to privilege one single vision above all others and to remain bound by it amidst a crowd of possible interpretations. Dots, colors, and shapes in an abstract painting may suddenly coalesce into a human face or a lunar landscape, wild geese fly out of a drawing, a monster stare at us from out the scrolls and patterns of a persian rug, and remain with us for ever.

Such a (literally) enthralling experience can likewise result from one’s encounter with a piece of fiction. This is notably so when the work has been the object of the multifarious scrutiny of academic exegesis over a relatively short span of time, which allows for the comprehensive yet comparative view. Interesting and enriching as all its interpretations may eventually prove to be, one of them mysteriously ravishes our imagination, never to set it free again.

Rising to the surface of the rich lode of commentary mined over the past thirty years or so since the novel’s revival, one such interpretation casts its spell over the figure of Edna Pontellier in Kate Chopin’s The Awakening. Indeed, since Sandra M. Gilbert’s flamboyant essay in The Kenyon Review,^{1} Edna in the scanty garb of Aphrodite is for ever swimming back to Cyprus or Cythera in the full glory of immortality… at least for the author of the present article who confesses to a susceptibility to mythology.

It can appear supererogatory to add yet another piece of exegesis to “one of the most persistently analyzed American novels,” to quote Gilbert’s own words. Yet Aphrodite beckons, and how could one resist her summons when gods and mortals alike, if the Homeric hymns are to be believed, were once powerless to do so ?

ANNE FOATA

This essay will take for granted the various stages of Edna Pontellier’s awakening and their discussion by the critics, all well known by now. What has been overlooked, however, is the emblematic nature of the event that befalls Edna during her fateful summer on Grand Isle. Edna actually experiences an epiphany, of the kind that from time immemorial has changed the lives of a happy (or unhappy) few : semi-celestials, mystics, artists, common men and women. She is granted an awareness of “her position in the universe as a human being ;”^{2} she is led to question the “realities” of her world, which sets her on a quest for “the abiding truth” (27). Edna, indeed, is made aware that “the realities pressing into her soul”

(32) have nothing to do with “the deeper undercurrents of life” (93) and with what she perceives to be her “real” self. The “real” woman within her clamors to be heard, recognized and fulfilled. During the following months of what can aptly be called the winter of her discontent, back in New Orleans, she keeps shedding her various societal personae like so many “pricking” garments. Our final glimpse of her finds Edna “free” at last and naked in the sun of the Grand Isle spring, about to swim back into her element and complete her rebirth as the triumphant goddess of love of Gilbert’s Kenyon Review article.

In dwelling upon the compelling figure of Aphrodite one should not forget that she is not one but two, a fact Gilbert failed to emphasize. And dual also is the “erotic” longing she fosters in the bodies and souls of men and women. Edna falls prey to the goddess of lustful love when she experiences the “first-felt throbbings of desire” (31) for Robert Lebrun, when her senses flame up under Alcée Arobin’s kiss (83), or when she proclaims herself a free woman, ready to “give [herself] where [she] choose[s]” (107). Edna, however, perceives the trap set by the senses and sensuality : a mere device engineered by Nature for the perpetuation of mankind, the illusion of the world.^{3} The abiding truth of which she is given a glimpse by Mademoiselle Reisz’ music must lie elsewhere, “beyond,” in a state of freedom which is harmony between herself and the universe, in the communion of her soul with the essence of life. Edna Pontellier, without ever so naming it (save to mention the “unattainable,” the “unlimited in which to lose herself”), is a seeker of the Absolute, absolute Love, absolute Beauty and Truth, which is the only Reality apt to fulfill the initiate who has been freed from the “mist,” the “life-long stupid dream” (107) of mortal existence.

APHRODITE REDUX

In other words, by awaking to the flesh, Edna sacrifices to the demands of Aphrodite Pandemos, the “common” or “popular” goddess of carnal desire, but her true yearning is for the blessings of Aphrodite Urania, the goddess of higher love which is also the all-inclusive Eros of Socrates’ encomium in Plato’s Symposium.^{4} The Pandemos and the Urania, the “common” and the “celestial,” refer to the two Aphrodites of Pausanias’ speech in the selfsame work.^{5}

To be sure, Pausanias’ eulogy of love must be taken with a grain of Attic salt ; after all, he is a Sophist, a member of a crew much derided by both Socrates and Plato for twisting reality for the sake of argument. His distinction between the two Aphrodites may not rest upon historical or mythological evidence ; it may even be alien to Plato’s philosophy proper ; it has nevertheless become part and parcel of our Western conception of love. As such, in their dichotomy between a higher and a lower form of love, the two Aphrodites epitomize Edna’s quest for the “significance of life” (83), which is also a search for her rightful position in the universe.

Edna’s assets in this respect are double, threefold even. The first is the common prerequisite, the sine qua non of any such endeavors : a penchant for inner contemplation and questioning, which enables her to set the illusion of her “actual” life against the demands of “real” life. The other two were recommended by Socrates and the wise woman of Mantinea who, he states in the Symposium, instructed him in the ways of Love. Like the seeker in Socrates’ lecture, Edna is not interested in securing immortality through the accumulation of generations. She may well be a hearty eater and drinker, a voluptuous sleeper and sensual lover, in a word, a splendid “animal,” she is also convinced, against the tide of accepted mores, that Fate has not fitted her to be a mother (20). Unwilling Pandemos, she is reluctant to contribute to the “grotesque pandemonium”

(58) of life, to add to the procession of humanity “struggling blindly [like worms] towards inevitable annihilation” (ibid). Weddings to her are among the “most lamentable spectacles on earth” (66). Her ultimate recoil from mere carnal love is triggered as much by her newly acquired awareness of the insatiability of desire as by a loathing for its ineluctable procreative consequences. Children eventually appear as “antagonists” in an alien world. All this is clear evidence that although she too occasionally sacrifices a goat to Aphrodite Pandemos, her secret longing is for the kind of love that makes men and women fecund according to the spirit and spells a higher form of self-fulfillment.

It is right that Edna should set off on her journey toward the ultimate truth of life endowed with that particular “susceptibility to beauty” (15) which is her third asset and the one that, according to Diotima, the Mantinean woman, is the first rung of the “ladder of love” leading to the “final revelation.” Chopin has Edna literally bathe in an atmosphere of beauty and sensuality, in which we see her “unfold[ing] like a torpid, torrid, sensitive blossom” (103).^{6} Much as it certainly owes to the wellknown fin de siècle thrill to which Chopin was certainly not immune, it is nevertheless the right atmosphere for an awakening to the higher spheres of Love, the ones transcending carnality, for as Diotima instructed Socrates who himself tells the revelers :

This is the way, the only way [to] the sanctuary of love. Starting from individual beauties, the quest for universal beauty must find him [the quester] ever mounting the heavenly ladder, stepping from rung to rung … to the special lore that pertains to nothing but the beautiful itself.. . . And remember, she said, that it is when he looks upon beauty’s visible presentment, and only then, that a man will be quickened with the true, not the seeming virtue … And when he has brought forth and reared this perfect virtue, he shall be called the friend of God, and if ever it is given to man to put on immortality, it shall be given to him (my emphasis).^{7}

Immortal, “unsullied, unalloyed, and freed from the mortal taint that haunts the frailer loveliness of flesh and blood,”^{8} such finally appears Aphrodite Urania, the winsome goddess of both life-force and universal wisdom Edna has been seeking through her mortal endeavors. And in her “handsomeness” and perfection of limb,^{9} such as Praxiteles once exalted them in the marble, also stands Edna Pontellier at the end of her trials in this our mortal world. She is about to yield once again to the “soft, close embrace” of the sea and, Aphrodite Anadyomene, to give herself back to the element out of which she was born. She is forever sailing back to Cythera.

Université Marc Bloch, Strasbourg, France.

NOTES

1 “The Second Coming of Aphrodite : Kate Chopin’s Fantasy of Desire,” The Kenyon Review, V (1983), 42–56.

2 p. 14–15 in the Norton Critical Edition, ed. by Margaret Culley (New York : W.W. Norton & Co., 1976). The pages of further quotations will hereafter appear parenthetically in the text of this article.

3 She also discovers that lust isn’t love, after Alcée Arobin’s first kiss. His kiss is a “flaming torch that kindled desire,” but she feels a “dull pang of regret because it was not the kiss of love which had inflamed her, because it was not love which had held this cup of life to her lips” (83).

4 The Collected Dialogues of Plato, Bollingen Series LXXI (New York : Pantheon Books, 1961).

5 A few words on the two goddesses’ background may answer the questions some readers must be raising at this point. Pausanias in his speech on Love during the banquet which is the motif of the Symposium mentions Aphrodite Urania as the older goddess, the one whose birth from the foam surrounding the castrated genitals of Uranus is described in Hesiod’s Theogony, whereas Aphrodite Pandemos is the daughter of Zeus and Dione, one of the three thousand Oceanids. Both Aphrodites have in truth co-existed from time immemorial and were not, originally at least, judgmentally opposed as the goddesses of higher and lower forms of Love as they appear in Pausanias’ speech. According to another Pausanias, the second century A.D. author of a Description of Greece, it was Aegeus, the legendary king of Athens, who introduced the cult of Aphrodite Urania to his city, on the model of the great oriental goddesses of the sky, the Sumerian Inanna or the Syrian Ishtar whom the Greeks renamed Astarte. Aegeus’ son, Theseus, then introduced the cult of Aphrodite Pandemos, “common” to all the Greeks, or at least to the peoples of the newly federated state of Attica. “Pandemos,” it is to be noted, had no ethical meaning originally ; it was a purely political or geographical determinative (Robert Graves in The Greek Myths translates “pandemos” by “federal,” making of Aphrodite Pandemos “Federal Aphrodite”). Both Aphrodites refer to older forms of deity, to the great mother-goddesses, the mistresses of life and death ; the lovely, frolicsome “golden” Aphrodite is an invention of the poets, Hesiod, Homer and others, followed by the sculptors and later the painters, who made her into the ideal of feminity and love, of the sexual urge whch spares neither gods, mortals or animals. (Robert Graves, The Greek Myths, Penguin Revised edition (Harmondsworth, 1960), see the index in vol. II, p. 380–81 for the multiple references to Aphrodite. My other source, besides Plato’s Symposium and Hesiod’s Theogony (in Penguin Classics, translated and edited by Dorothea Wender (Harmondsworth, 1973), lines 189–210) was Erika Simon : Die Götter der Griechen (Munich, Hirmerverlag, 1969), p. 229–254 ; 329–330).

6 or to shift the simile to the animal world, like “some beautiful sleek animal waking up in the sun,” “palpitant with the forces of life” (70), as she appears to old Dr Mandelet. “A subtle current of desire passed through her body,” (58) Chopin writes, when Edna thinks of Robert Lebrun ; the touch of Victor Lebrun’s lips upon her palm during the farewell dinner of chapter 30 is “like a pleasing sting on her hand” (90). Alcée Arobin’s experienced caresses make her “supple to his gentle seductive entreaties” (93) ; his kiss is “a flaming torch that kindled desire” (83). We see Edna also respond to the “soft physical contact,” the “excessive physical charm” of the sensuous Creole Madonna, Adèle Ratignolle (15–16), and submit to Mademoiselle Reisz’ ambiguous kiss (on her bare shoulder) and banter (“Bonne nuit, ma reine : soyez sage” (88)) on her leaving the farewell banquet. It is Mademoiselle Reisz’ playing Chopin’s Impromptus and Wagner’s Liebestod that gives Edna her “first impress of the abiding truth” (27). Edna’s first swim in the Gulf, with her body enclosed in the “soft, close embrace” (15) of the sea and the “quick vision of death [smiting] her soul” (29) rings an unmistakably orgiastic or orgasmic note. And last but not least, to add to the sensuousness of the whole atmosphere of the novel, there is the lush landscape of sub-tropical Southern Louisiana with its sun and breeze, its citrus trees and flowers, its odors and colors.

7 translated by Michael Joyce for The Collected Dialogues of Plato in the Bollingen series and quoted by Joseph Campbell : The Masks of God : Occidental Mythology (New York : The Viking Press, 1964), reprinted by Penguin Books (Harmondsworth, 1976), p. 231–232.

8 Ibid, p. 232.

9 cf. her portrait in chapters II (5) and VII (16).